# The Role of Integrity Monitoring in Connected and Automated Vehicles: Current State-of-Practice and Future Directions

### v1

@@ III-A In-vehicle (Standalone) Integrity Monitoring
III-B Cooperative Integrity Monitoring
@@
In addition to the standard parameters used in the traditional IM framework, i.e. accuracy, availability, continuity, integrity, two other parameters - timeliness & interoperability, have been proposed in the literature to evaluate the performance of cooperative-ITS (C-ITS) applications [37]. In [38], a runtime IM framework for real-time relative positioning was introduced focusing on sharing raw GPS observation data through Dedicated Short Range Communications (DSRC). The author of previous work also explored a real-time relative position prediction strategy [39] in a V2V environment using several position prediction techniques to compensate for the loss of the V2V DSRC link. In this work, only road-level accuracy (<<<5 m error) could be guaranteed. A hybrid integrity monitoring method is proposed in [40] that overcomes the limitations of conventional RAIM technique under limited GNSS satellite coverage. It generates virtual satellite measurements using DSRC range rate measurements from neighboring vehicles and road map data. Recently, signals of opportunity (SOPs) such as AM/FM radio, cellular signals, etc., have aided the IMU in the absence of GNSS signals. Cellular signals in particular are of great interest due to their availability in urban canyons. A SOP-based RAIM framework was developed in [41] which fuses cellular pseduorange measurements with IMU measurements. The proposed method showed a reduction of position Root Mean Square Error (RMSE) by 66%.
@@
Paper Authors Year Mode Data Source Techniques
[17] Hassan et al. 2021 Standalone GNSS, IMU, Odometry, LiDAR RAIM, KF-based
[25] Worner et al. 2016 Standalone N.A FDI, BE approach
[27] Jing et al. 2022 Standalone & Cooperative GNSS/IMU, Map, Wireless RAIM, KF-based
[22] Raouf et al. 2022 Standalone GNSS, Camera, LiDAR, Radar RAIM, KF-based
[20, 30] Balakrishnan et al. 2020 Standalone GPS, Camera, LiDAR, Map RAIM, KF-based
[28, 29] Gottschalg et al. 2021 Standalone GNSS, IMU, Odometry kSigma, KIPL, ARAIM
[38] Ansari et al. 2015 Cooperative GNSS/RTK, IMU, DSRC PRMF, RAIM, KF-based fault detection
[39] Ansari 2020 Cooperative GNSS, DSRC KF, EKF, PF, MR, MFNN
[40] Liu et al. 2020 Cooperative GNSS, DSRC, Road map Hybrid-RAIM, Virtual satellite measurements
[41] Maaref et al. 2022 Cooperative GNSS, IMU, Cellular-LTE SOP-based RAIM
[45] Xiong et al. 2021 Cooperative GNSS, UWB CIM, Kalman residual
[47] Schon et al. 2018 Cooperative GNSS, IMU, stereo camera, laser scanner Interval set theory
The development of cooperative integrity monitoring methods is still in its early stages. The question remains whether the empirically determined Required Navigation Performance (RNP) values hold consistent across different driving conditions, positioning hardware, and communication topologies. Inconsistencies in RNP parameters in CAV applications would lead to multiple solutions from various automotive OEMs resulting in loss of standardization. In [48], the authors have defined the minimum operational performance standards that can be adjusted to meet the needs of a specific safety application based on the positioning capabilities of a vehicle. The ESA-funded P-CAR project [49] focuses on identifying and validating connected and automated driving (CAD) functions to meet safety integrity needs. It aims at creating realistic Hardware-in-Loop (HiL) simulations and appropriate fault injection strategies to evaluate events with integrity risks that might be rare and difficult to test in real-world conditions. Based on the initial stakeholder surveys, truck platooning is being favored to test the efficacy of V2V communication, whereas, applications such as highway pilot, urban automated shuttle, and valet parking are key contenders in the navigation segment. Various cooperative integrity monitoring strategies and key findings are summarized in TABLE II.
@@ IV Automotive Safety Standards and Protocols
TABLE IV: Summary of C-ITS Projects and Deployments In Europe
NordicWay [78]
- NordicWay 1
- NordicWay 2
- NordicWay 3
EU, VOLVO,
SAFEROAD, SKANSKA, BM SYSTEM, etc.
- 2015-2017
- 2017-2020
- 2019-2023
@@
TABLE V: Summary of C-ITS Projects and Deployments In USA
USDOT Connected Vehicle Pilots [102]
- New york city pilot
- Tampa (THEA) pilot
- Wyoming DOT pilot
@@ V Real-world Cooperative Driving Automation Experiments and Public-Datasets
These V2X datasets mostly focus on cooperative perception and sensor fusion problems and offer a benchmark for 3D detection, tracking, and trajectory prediction analyses. However, the dataset is collected locally at source i.e. vehicle, infrastructure, and doesn’t involve active V2X communication among the agents. The communication constraints such as bandwidth, latency, packet drops, etc., are imposed in post-processing analysis. Recently released datasets such as Berlin-V2X, V2AIX, and TiHAN-V2X, involve active communication among the agents and have collected data for evaluating real-time communication performance. Berlin-V2X [111] and TIHAN-V2X [112] primarily provide information on cellular and sidelink communication in V2X scenarios that can help evaluate Quality of Service (QoS) Prediction and ML algorithms focusing on optimizing the communication channel. V2AIX dataset [113], released in 2024, consists of both cooperative perception data (V2V & V2I) and ETSI ITS-G5 V2X messages such as CAM, DENM, MAPEM, and SPATEM. Such datasets are crucial for analyzing the effects of CAV penetration rate on various applications under real-world communication constraints.
@@ Acknowledgments
[Uncaptioned image] Saswat Priyadarshi Nayak received the B. Tech degree in Electrical Engineering from the National Institute of Technology Rourkela, India in 2018. He served as a Project Associate at the Department of Aerospace Engineering, Indian Institute of Technology Kanpur, India 2018-19. He is currently pursuing a Ph.D. degree at the Center of Environmental Research and Technology (CE-CERT), University of California Riverside, USA. He is one of the recipients of the 2024 ITS California (ITS CA) and California Transportation Foundation (CTF) Scholarship. His main research interests include vehicle positioning and localization in mixed traffic scenarios, multi-sensor fusion, and connected vehicle applications.
[Uncaptioned image] Matthew J. Barth (Fellow, IEEE) received the M.S. and Ph.D degree in electrical and computer engineering from the University of California at Santa Barbara, in 1985 and 1990, respectively. He is currently the Hays Families Professor in the College of Engineering, University of California at Riverside, USA. He also served as the Director of the Center for Environmental Research and Technology. His current research interests include ITS and the environment, transportation/emissions modeling, vehicle activity analysis, advanced navigation techniques, electric vehicle technology, and advanced sensing and control. Dr. Barth has been active in the IEEE Intelligent Transportation System Society for many years, serving as a Senior Editor for both the Transactions of ITS and the Transactions on Intelligent Vehicles. He served as the IEEE ITSS President for 2014 and 2015 and is currently the IEEE ITSS Vice President of Education.

### v2

@@ III-B In-vehicle (Standalone) Integrity Monitoring
Paper Year Mode Data Source Techniques Key Takeaways (+) & Identified Gaps (-)
Hassan et al. [10] 2021 Standalone GNSS, IMU, Odometry, LiDAR Least-squares-based ARAIM (+) : Suitable for loosely coupled integration with other sensors and multi-faul multi-constellation scenarios
(-) : Need for IM architectures for diverse non-GNSS sensors or positioning solutions
Worner et al. [29] 2016 Standalone N.A BE - Interval analysis (+) : Guaranteed bound on PE, Robust to outliers
(-) : Conservative, Computationally intensive, scalability issue for large state space
Jing et al. [8] 2022 Standalone & Cooperative GNSS/IMU, Map, Wireless RAIM (+) : Classical RAIM is easy to implement and suitable for real-time operations
(-) : Lack of comparable protection level (PL) computation from existing IM methods under the same vehicle testing environment
Raouf et al. [26] 2022 Standalone GNSS, Camera, LiDAR, Radar KF-based RAIM (+) : Fault identification and diagnosis for overall sensors of ADAS for AVs
(-) : Need for comprehensive prognostic health management considering variation in vehicle type, sensor modality & algorithms
Balakrishnan et al. [24, 36] 2020 Standalone GPS, Camera, LiDAR, Map Model/Coherence-based IM (+) : Robust against perception errors- partial detections, occlusions
(-) : Inability to detect common features in multi-modal sensor data
Gottschalg et al. [33, 40] 2021 Standalone GNSS, IMU, Odometry kSigma, KIPL, ARAIM (+) : KIPL outperforms other methods in terms of accuracy and availability
(-) : Integrity algorithms yield varying protection levels, affecting availability and requiring selection based on application
Ansari et al. [34] 2015 Cooperative GNSS/RTK, IMU, DSRC PRMF, RAIM, KF-based fault detection (+) : Real-time IM benchmarking with DSRC availability monitoring
(-) : BSM reception affected by crowded/saturated network condition
Ansari [46] 2020 Cooperative GNSS, DSRC KF, EKF, PF, MR, MFNN (+) : Motion Model and Maneuver Recognition (MR) outperformed other methods in accuracy and computation complexity
(-) : Discussed motion prediction models lack precision for advanced safety tasks
Liu et al. [47] 2020 Cooperative GNSS, DSRC, Road map Hybrid-RAIM, Virtual satellite measurements (+) : Overcomes limited GPS coverage constraints under RAIM unavailability
(-) : Doesn’t outperform conventional RAIM in open-sky condition
Maaref et al. [48] 2022 Cooperative GNSS, IMU, Cellular-LTE SOP-based RAIM (+) : Cellular pseudoranges aiding for IMU in the absence of GNSS signals (Urban canyons)
(-) : Prior knowledge of cellular tower position, more multipath error compared to GNSS, estimation of cell-tower clock bias
Xiong et al. [49] 2021 Cooperative GNSS, UWB Kalman residual-based CIM (+) : High fault detection sensitivity compared to RAIM, fault detection and diagnosis in cooperative positioning
(-) : UWB ranging affected by Non-Line-of-Sight (NLOS) errors, Needs multi-sensor testing and validation
Schon et al. [50] 2018 Cooperative GNSS, IMU, stereo camera, laser scanner Interval set theory (+) : Boundedness of estimates & time-synchronization error in multi-sensor system
(-) : Overestimation of error bounds
III-C Cooperative Integrity Monitoring
@@
In addition to the standard parameters used in the traditional IM framework, i.e. accuracy, availability, continuity, integrity, two other parameters - timeliness & interoperability, have been proposed in the literature to evaluate the performance of cooperative-ITS (C-ITS) applications [51]. In [34], a runtime IM framework for real-time relative positioning was introduced focusing on sharing raw GPS observation data through Dedicated Short Range Communications (DSRC). The author of previous work also explored a real-time relative position prediction strategy [46] in a V2V environment using several position prediction techniques to compensate for the loss of the V2V DSRC link. In this work, only road-level accuracy ($<$5 m error) could be guaranteed. A hybrid integrity monitoring method is proposed in [47] that overcomes the limitations of conventional RAIM technique under limited GNSS satellite coverage. It generates virtual satellite measurements using DSRC range rate measurements from neighboring vehicles and road map data. Recently, signals of opportunity (SOPs) such as AM/FM radio, cellular signals, etc., have aided the IMU in the absence of GNSS signals. Cellular signals in particular are of great interest due to their availability in urban canyons. A SOP-based RAIM framework was developed in [48] which fuses cellular pseduorange measurements with IMU measurements. The proposed method showed a reduction of position Root Mean Square Error (RMSE) by 66%.
@@
The development of cooperative integrity monitoring methods is still in its early stages. The question remains whether the empirically determined Required Navigation Performance (RNP) values hold consistent across different driving conditions, positioning hardware, and communication topologies. Inconsistencies in RNP parameters in CAV applications would lead to multiple solutions from various automotive OEMs resulting in loss of standardization. In [56], the authors have defined the minimum operational performance standards that can be adjusted to meet the needs of a specific safety application based on the positioning capabilities of a vehicle. The ESA-funded P-CAR project [57] focuses on identifying and validating connected and automated driving (CAD) functions to meet safety integrity needs. It aims at creating realistic Hardware-in-Loop (HiL) simulations and appropriate fault injection strategies to evaluate events with integrity risks that might be rare and difficult to test in real-world conditions. Based on the initial stakeholder surveys, truck platooning is being favored to test the efficacy of V2V communication, whereas, applications such as highway pilot, urban automated shuttle, and valet parking are key contenders in the navigation segment. Various cooperative integrity monitoring strategies and key findings are summarized in TABLE II. While aspects such as computational demand and latency are important for deployment, a quantitative, apples-to-apples comparison would require a dedicated meta-analysis or controlled simulation studies across diverse positioning techniques, sensor modalities, cooperation level (V2V, V2I, etc.), and environmental conditions (open-sky, highway, tunnel), which is beyond the scope of this paper.
@@ IV Automotive Safety Standards and Protocols
TABLE IV: Summary of C-ITS Projects and Deployments In Europe
NordicWay [77]
- NordicWay 1
- NordicWay 2
- NordicWay 3
EU, VOLVO,
SAFEROAD, SKANSKA, BM SYSTEM, etc.
- 2015-2017
- 2017-2020
- 2019-2023
@@
TABLE V: Summary of C-ITS Projects and Deployments In USA
USDOT Connected Vehicle Pilots [93]
- New york city pilot
- Tampa (THEA) pilot
- Wyoming DOT pilot
@@
Figure 6: Timeline of major Cooperative ITS and Connected Vehicle pilot projects in Europe (top row) and the United States (bottom row) from 1999 through 2025. Projects marked with an asterisk (*) are currently active.
@@ V Real-world Cooperative Driving Automation Experiments and Public-Datasets
These V2X datasets mostly focus on cooperative perception and sensor fusion problems and offer a benchmark for 3D detection, tracking, and trajectory prediction analyses. However, the dataset is collected locally at source i.e. vehicle, infrastructure, and doesn’t involve active V2X communication among the agents. The communication constraints such as bandwidth, latency, packet drops, etc., are imposed in post-processing analysis. Recently released datasets such as Berlin-V2X, V2AIX, and TiHAN-V2X, involve active communication among the agents and have collected data for evaluating real-time communication performance, as shown in fig7. Berlin-V2X [119] and TIHAN-V2X [122] primarily provide information on cellular and sidelink communication in V2X scenarios that can help evaluate Quality of Service (QoS) Prediction and ML algorithms focusing on optimizing the communication channel. V2AIX dataset [121], released in 2024, consists of both cooperative perception data (V2V & V2I) and ETSI ITS-G5 V2X messages such as CAM, DENM, MAPEM, and SPATEM. Such datasets are crucial for analyzing the effects of CAV penetration rate on various applications under real-world communication constraints.
Figure 7: Evolution of V2X datasets from object detection & tracking to segmentation & trajectory forecasting, with a growing number of recent releases emphasizing CAV communication performance.